OpenStudy (anonymous):

When one U.S. Dollar is worth one point seven four Australian Dollars, which of the following could be used as a conversion factor? five hundred eighty-sevenths one and sixteen twenty-fifths one and thirty-seven fiftieths two and thirteen fiftieths

OpenStudy (anonymous):

In fact, 1 and 2/5 = 1+2/5 = (5+2)/5 = 7/5 which is equal to 1.4

OpenStudy (anonymous):

i needd it in a conversation factor plzzz